### v1

# QuickVideo: Real-Time Long Video Understanding with System Algorithm Co-Design

Abstract
Long video understanding has emerged as a crucial capability in real-world applications such as meeting summarization, video surveillance, educational lecture analysis, and content moderation. However, it remains computationally prohibitive for VideoLLMs, primarily due to two bottlenecks: 1) sequential video decoding, the process of converting the raw bit stream to RGB frames can take up to a minute for hour-long video inputs, and 2) costly prefilling of up to several million tokens for LLM inference, resulting in high latency and memory use. To address these challenges, we propose QuickVideo, a system-algorithm co-design that substantially accelerates long video understanding to support real-time downstream applications. It comprises three key innovations: QuickCodec, a parallelized CPU-based video decoder that achieves 2–3$\times$ speedup by splitting videos into keyframe-aligned intervals processed concurrently. QuickPrefill, a memory-efficient prefilling method using KV-cache pruning to support more frames with less GPU memory; and an overlapping scheme that overlaps CPU video decoding with GPU inference. Together, these components reduce the time required to process a long video input by a minute, enabling fast, efficient video understanding even on limited hardware. Experiments show that QuickVideo generalizes across durations and sampling rates, making long video processing feasible in practice.
1 Introduction
Video data has become the dominant modality for conveying information online. As of 2023, video data accounts for two thirds of all data transmitted over the Internet [30]. Much of this data is “long video” ranging from minutes to hours in duration, from online conferencing, gaming, social networking, and movie streaming. This torrent of online video data demands efficient and automated understanding for problems such as content moderation [2], real-time surveillance [43], and accessibility [22]. Video Large Language Models (VideoLLMs) [4, 48, 7] have emerged as powerful tools to support these downstream tasks. By natively processing entire video inputs, VideoLLMs exhibit phenomenal potential to understand and reason about video content, offering a practical solution for managing and extracting information from the exponentially growing flood of video data across the Internet [47].
However, using VideoLLMs for long video understanding suffers from several efficiency challenges. First, the entire video must be decoded from raw bitstreams into RGB frames before the model can begin processing. Current frameworks require up to a minute to decode the frames from an hour-long video input, introducing high latency before any context prefill can start. Second, the prefilling step itself is both computationally and memory intensive [37]. Each frame—representing an instantaneous moment—can consume hundreds of tokens in the model context [49, 8]. As a result, even a modest frame rate (e.g., 2 FPS) for an hour-long video can lead to millions of tokens, far exceeding the memory budget of standard GPUs. Qwen2.5-VL [4] introduced several architecture modification to accelerate video processing. However, using Qwen2.5-VL-7B, prefilling an hour-long HD video sampled at its native $2$ fps still requires more than the $80$ GB of memory offered by an A100/H100 SXM4 GPU. Even after reducing the frame sampling rate by 4$\times$, prefilling still takes over 25 seconds on datacenter-grade hardware. These inefficiencies result in a frustrating user experience, characterized by long delays and prohibitively high hardware requirements. Users with limited computational resources are effectively excluded from accessing the long video understanding capabilities of VideoLLMs.
Figure 1: An overview of how QuickVideo overlaps video decoding on CPU (QuickCodec) and prefill on GPU (QuickPrefill). QuickCodec concurrently processes intervals of the compressed video bit stream. QuickPrefill uses independent groups of frames, therefore it can begin prefill once the first frames are decoded, outputting carefully selected KV vectors. As QuickCodec loads frames synchronously, QuickPrefill can process the next prefill group immediately. This results in video decoding and prefill being almost enirely overlapped.
To mitigate the computational overhead of long video VideoLLMs use extremely low frame sampling rates when processing long video inputs, instead of their native 1-2 FPS [4, 7]. Frames are sampled as much as a minute apart during hour-long video understanding [49, 8]. A minute gap between sampled frames can result in missing crucial video segments required for an understanding task. Low frame sampling rates also make fine-grained temporal and motion understanding impossible, as intervening frames are mostly removed [25]. Effective long video understanding thus requires loading and prefilling thousands of frames while preserving temporal continuity. Developing faster, more efficient VideoLLMs is critical for enabling comprehension of videos that span hours.
Currently, video decoding and context prefilling are treated as disjoint and sequential stages in the VideoLLM pipeline. Moreover, video decoding is largely overlooked, despite contributing substantially to end-to-end latency. To remedy this, we introduce QuickVideo, a framework for faster, memory-efficient long video understanding. QuickVideo reduces the latency and resource requirements of these key bottlenecks in long video understanding. Our framework empowers fast video understanding on video inputs consisting of hundreds of thousands of frames, while maintaining the sampling rates required for fine-grained understanding. QuickVideo introduces three core contributions for accelerating long video understanding in VideoLLMs:
(1) System-Level $\rightarrow$ QuickCodec: a drop-in replacement video decoder designed for VideoLLMs. By redesigning video decoding for VideoLLM frame sampling, we achieve a 2-3x speedup compared to existing libraries when loading hour-long video inputs.
(2) Algorithm-Level $\rightarrow$ QuickPrefill: a group-based prefilling strategy combined with key-value (KV) cache pruning, which significantly reduces both computation time and memory usage during the prefilling stage, while incurring less than 3% accuracy degradation in most benchmarks.
(3) Co-Design $\rightarrow$ Overlapped Execution Scheme: the strategy tightly couples CPU-based QuickCodec and GPU-based QuickPrefill, enabling near-complete overlap to maximize efficiency. QuickVideo reduces the time to infer a 30 minute video input by more than 3x, from $69.7$ seconds to only $20.0$ seconds. The results demonstrate the effectiveness of our system-algorithm co-design.
2 Background
We provide an overview of VideoLLM inference and key concepts in video processing. Although details vary, this background is broadly applicable to standard VideoLLM architectures and video standards. For clarity, we use “video decoding” to describe the process of decoding the compressed video into a tensor of video frames, and use “LLM decoding” to denote the process of auto-regressive decoding of a large language model.
2.1 VideoLLM Inference
VideoLLMs must first decode a compressed video into a packed frame tensor before tokenization. The resulting raw frames are then passed through a visual encoder, which converts them into video tokens suitable for input to the LLM. Unlike text preprocessing, which relies on lightweight tokenizers, video decoding is inherently slow on both CPU and GPU due to its sequential nature [38, 31]. Despite this, prior work in LLM video understanding has largely overlooked the latency incurred by this stage. Following preprocessing, the generation process of a VideoLLM consists of two stages: (1) Prefill, where both video and text tokens are processed to compute key-value (KV) caches for each transformer layer; and (2) LLM decoding, where tokens are generated autoregressively using the stored KV representations. The prefill stage is computationally expensive due to the quadratic complexity $\mathcal{O}(n^{2})$ of self-attention over long sequences, while the decoding stage is memory-intensive as it requires storing and repeatedly accessing the full KV cache.
Let $\mathbf{X}^{v}=\{\mathbf{x}_{1}^{v},\ldots,\mathbf{x}_{|\mathbf{X}^{v}|}^{v}\}$ and $\mathbf{X}^{t}=\{\mathbf{x}_{1}^{t},\ldots,\mathbf{x}_{|\mathbf{X}^{t}|}^{t}\}$ represent the video and text tokens, respectively, with video tokens preceding the text. For each transformer layer $l\in\{1,\ldots,L\}$, the KV cache comprises tensors $\mathbf{K}^{(l)},\mathbf{V}^{(l)}\in\mathbb{R}^{(|\mathbf{X}^{v}|+|\mathbf{X}^{t}|)\times n_{h}\times d_{h}}$, where $n_{h}$ is the number of attention heads and $d_{h}$ is the per-head dimensionality. For example, let 8B InternVL-2.5 [7] model process a one-hour video at 1 frame per second, the total required memory is around 400GB (see subsection D.2). This memory footprint makes KV cache storage a critical bottleneck in VideoLLM inference, significantly limiting the maximum processable video length and constraining the feasible batch size.
2.2 Long Video Processing
Multimedia container formats like MP4 or MKV bundle all the elements required for media playback, including video streams, audio streams, subtitles, and metadata [18]. In these containers, videos are stored as compressed bit streams [18, 38]. In multimedia processing libraries like FFmpeg [32], video decoding is described by a queue $\mathcal{D}$ that enqueues fixed-sized blocks of the bit stream, called packets, as input and dequeues video frames. We denote a bit stream $\mathcal{S}=(p_{0},p_{1},\dots,p_{n-1})$ and a video $\mathcal{V}=(f_{0},f_{1},\dots,f_{m-1})$ as ordered lists of packets and frames, respectively. Each frame $f_{i}$ is a tensor containing 8-bit integers of shape $(3\times h\times w)$, where $h$ is the pixel height and $w$ is the pixel width. In general, packets are not frame aligned, enqueueing a single packet to the decoder can cause the decoder to output zero, one or potentially multiple frames [38]. This is because frames require varying amounts of information to encode, and therefore cannot be aligned to fixed-sized packets. Furthermore, video frames are not encoded independently in bit stream, as surrounding frames contain redundant information. Therefore, the video encoder encodes the residual of the frame in the bitstream, instead of the frame itself [^1] The encoded residual of a frame may require information from previous or future frames to decode [38]. [38, 31]. For this reason, video decoding is a largely sequential process, where previous frames must decoded first and then the residual information encoded in the bit stream can be used to decode the next frame [38]. Although the video encoder may also reorder frames in the bit stream for efficiency, the decoder always outputs frames in the order that they should be displayed during playback [32].
Packet and Frame Metadata. Although metadata is not directly encoded in the bit stream or frame itself, for simplicity, we denote metadata corresponding to packets or frames as if they are fields. The packet and frame metadata is stored in the container, not the bit stream [18]. The presentation timestamp (pts) of a frame is a 64-bit unsigned integer that represents when the frame should be displayed to a user [32]. Most formats do not include global frame positioning information in metadata. We instead use Equation 1 to rescale the presentation timestamp for a frame $f$ to obtain $f$’s index $i$ in $\mathcal{V}$.
$i=\left\lfloor\frac{(m-1)\cdot f\text{.pts}}{pts_{max}-pts_{min}}\right\rceil$ (1)
$pts_{max}$ and $pts_{min}$ are the minimum and maximum presentation timestamp for the video stream. Each packet has a keyframe flag that marks that video decoding can begin from its position [18, 32].
2.3 Keyframes and Seeking
As video decoding relies on surrounding frames, it is a sequential process. However, during playback, users may want to navigate and skip through the video. To support this, the bit stream contains keyframes, which act as reset points from which video decoding can begin. Keyframes are encoded at semi-regular intervals in $\mathcal{S}$, usually a few seconds apart. To use keyframes to navigate in $\mathcal{S}$, we use the Seek subroutine. Seek(${\mathcal{S},\;pts}$) finds the keyframe packet $p_{i}\in\mathcal{S}$ such that decoding from $p_{i}$ yields all $f$ such that $f\text{.pts}\geq pts$. However, seeking introduces overhead, as it requires flushing decoder buffers and reinitializing state [32].
Algorithm 1 Seek-based video decoding
1:Bit stream $\mathcal{S}$, Ordered set $\mathcal{I}$, Video Decoder $\mathcal{D}$, $h$, $w$
2:Allocate memory block ${\bm{F}}$ of size ||×$|\mathcal{I}|\times 3\times h\times w$
3:for $i\in\mathcal{I}$ do
4:     Estimate $pts$ of $f_{i}$
5:     $p_{i}\leftarrow\textsc{Seek}(\mathcal{S},\;pts)$triangleright Seek to the keyframe before $f_{i}$ in $\mathcal{S}$
6:     Decode $p_{i},p_{i+1},\dots$ until $\mathcal{D}$ outputs $f_{i}$
7:     Write $f_{i}$ to ${\bm{F}}$
8:return ${\bm{F}}$
Algorithm 1 is a standard approach when decoding video for machine learning [12, 27]. For each desired frame $f_{i}$, given by selected indices in $\mathcal{I}\subseteq\{1,2,\dots,m-1\}$, the algorithm does the following: It seeks for the keyframe closest to $f_{i}$ in $\mathcal{S}$, and then it decodes packets until $\mathcal{D}$ outputs $f_{i}$. $f_{i}$ is saved in the buffer ${\bm{F}}$. This algorithm performs well for sparse access patterns, as if there are large gaps between desired frames, seeking before decoding each frame is ideal.
3 Method
In this section, we first introduce QuickVideo, mainly consisting of three components:
3.1 QuickCodec: Long Video Decoding for VideoLLMs
Given a bit stream $\mathcal{S}$ for a video $\mathcal{V}=(f_{1},f_{2},\dots,f_{m})$ with frame height $h$, frame width $w$ and a desired degree of concurrency $c$, our goal is to compute ${\bm{F}}$ such that $\forall j\in\{0,1,\dots,|\mathcal{I}|-1\}\;{\bm{F}}_{j}=f_{\mathcal{I}[j]}$ for $\mathcal{I}\subseteq\{1,2,\dots,m-1\}$. That is, ${\bm{F}}$ is a packed tensor containing all the frames selected in $\mathcal{I}$. We assume that $m$ is known from container metadata or an estimate using $pts_{max}$ and $pts_{min}$. The efficiency of our algorithm relies on two observations:
(1) It is faster to use $c$ cores to decode $c$ short videos than use $c$ cores to sequentially decode $1$ long video. Video decoding for human playback focuses on the second case, as humans watch earlier frames while later frames decode. However, due to interframe dependencies, sequential video decoding is difficult to parallelize [38]. Unlike in human playback, VideoLLMs require the entire video input to be loaded upfront. Therefore, we can decompose the loading of a long video $\mathcal{V}$ into loading $c$ short videos that span $\mathcal{V}$. However, we cannot begin video decoding from an arbitrary frame, only a keyframe. Keyframe Intervals (Appendix A) is a subroutine that parses the metadata of $\mathcal{S}$ and computes $c$ approximately equal length intervals, starting and ending on keyframes, that span $\mathcal{V}$. We parallelize over these intervals in Algorithm 2.
(2) Ideally, VideoLLMs sample frames at a short, regular interval, usually 1-2 FPS [4]. This is less than the gap between keyframes in standard codecs. Therefore, seek-based decoding must decode from all keyframes regardless, and many seeks are redundant. Our algorithm only requires $1$ seek operation per core, instead of seeks proportional to the number of frames sampled.
Algorithm 2 QuickCodec
1:Bit stream $\mathcal{S}$, Ordered set $\mathcal{I}$, Video Decoder $\mathcal{D}$, $h$, $w$, $c$, $m$
[…]
3:Allocate shared memory ${\bm{F}}$ of size ||×$|\mathcal{I}|\times 3\times h\times w$
4:Initialize memory offset map $M$
5:for $k\in\{0,1,\dots,|\mathcal{I}|-1\}$ do
6:     $M[\mathcal{I}[k]]\leftarrow k$triangleright Maps frame index to memory offset in ${\bm{F}}$
7:for all  $(pts_{start},\;pts_{end})\in\mathcal{J}$ in parallel do $\triangleright$ Parallelize over $t$ intervals
8:     $p_{i}\leftarrow\textsc{Seek}(\mathcal{S},\;pts_{start})$triangleright Seek to the packet at the start of the keyframe interval
9:     repeat
10:         while $\mathcal{D}$ not empty do
[…]
12:              if $f\text{.pts}\geq pts_{end}$ then
13:                  break
14:              Compute $i$ with equation 1
15:              if $i$ in $M$ then
16:                  $o\leftarrow M[i]$triangleright Get the memory offset for $f_{i}$ in ${\bm{F}}$
17:                  ${\bm{F}}_{o}\leftarrow f$triangleright Write frame into shared memory tensor
18:         $\mathcal{D}\text{.enqueue($p_{i}$)}$
19:         $p_{i}\leftarrow p_{i+1}$triangleright Get next packet in bit stream $\mathcal{S}$
20:     until $f\text{.pts}\geq pts_{end}$
21:return ${\bm{F}}$
Algorithm 2 describes the core of our video decoding algorithm. The algorithm begins by using metadata to calculate $c$ keyframe-aligned intervals span the video (line 1). Lines 2-5 initialize a block of shared memory ${\bm{F}}$ and compute a dictionary $M$ that maps indices of selected frames to unique memory offsets in ${\bm{F}}$. We then decode the long video in $c$ parallel intervals (lines 6-19). Video decoding begins by seeking the the start of the interval $pts_{start}$, which is guaranteed to be a keyframe (line 7). We enqueue packet to decode (lines 17-18) until the video decoder yields frames to process (line 9). If the timestamp of the dequeued frame is greater than or equal to the endpoint of the interval $pts_{end}$, parallel processing ends (lines 11-12 and 19). As the intervals in $\mathcal{J}$ span $\mathcal{S}$, $pts_{min}$ and $pts_{max}$ are given by the least and greatest values in $\mathcal{J}$, respectively. Therefore, we use can Equation 1 to calculate the index $i$ of $f$ (line 13). Lastly, we save $f$ to ${\bm{F}}$ if $f$ is a selected frame (lines 14-16). As decoding from a keyframe yields all frames with greater $pts$ and $\mathcal{D}$ outputs frames in $pts$ order, when the parallelized loop exits (line 19), all selected frames with $pts$ in the interval $[pts_{start},pts_{end})$ will have been output by $\mathcal{D}$ and saved to ${\bm{F}}$. Therefore, as $\mathcal{J}$ spans $\mathcal{S}$, when the algorithm returns, ${\bm{F}}$ will contain all selected frames.
3.2 QuickPrefill: Efficient Group-based Pre-filling for VideoLLM
After decoding the video bit stream into packed tensors, they are then fed into the VideoLLM for inference. However, LLM generation with a long context is a well-known challenge due to its intensive memory usage and costly computation. To mitigate this issue, we introduce QuickPrefill, a grouped prefilling and KV cache pruning method that both accelerates the speed and reduces the memory requirement significantly.
Group-based Prefill.
Let $V={\mathbf{v}_{1},\mathbf{v}2,\ldots,\mathbf{v}{|\mathbf{X}^{v}|}}$ denote the sequence of video tokens, where $|\mathbf{X}^{v}|$ is the total number of tokens and each token $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional vector. To reduce memory overhead during prefilling, we adopt a group-based strategy by partitioning the video token sequence into $G$ disjoint groups: $\mathbf{V}={\mathbf{V}_{1},\ldots,\mathbf{V}_{G}}$, where each group $\mathbf{V}_{g}$ contains approximately $N_{g}=\frac{|\mathbf{X}^{v}|}{G}$ tokens. Instead of processing the entire sequence at once, we sequentially prefill each group and store its corresponding key-value (KV) cache as $\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}$ for each transformer layer $l$. This strategy significantly reduces the peak activation memory usage by a factor of $G$ and remains effective even when used in conjunction with efficient attention mechanisms such as FlashAttention. Empirically, it enables hour-long video understanding while keeping GPU memory usage within practical limits (e.g., reducing more than 100 GB of memory overhead; see subsection D.2).
Group-based KV Cache Pruning.
While group-based prefill can effectively reduce the peak activation memory, the largest memory bottleneck, KV cache memory is still not resolved. Therefore, when processing each group of video tokens, instead of saving the all KV cache vectors, we will prune unimportant KV cache to maintain an $\rho\in(0,1]$ retention ratio and thus reduce the KV cache memory usage by a factor of $\frac{1}{\rho}$. The pruning decision is based on an importance score function $\mathbf{s}$ to produce an ordered list where we select the top-k KV cache until the retention ratio is reached:
$\tilde{\mathbf{K}}_{g}^{(l)}=\mathbf{K}_{g}^{(l)}[I_{g}^{(l)}],\text{ }\tilde{\mathbf{V}}_{g}^{(l)}=\mathbf{V}_{g}^{(l)}[I_{g}^{(l)}],\text{ where }I_{g}^{(l)}=\operatorname{TopK}\left(\mathbf{s}(\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}),\;k=\rho\cdot N_{g}\right)$ (2)
where $\operatorname{TopK}$ defines a function that returns of the indices the top-k largest inputs. There are multiple heuristic importance score function $\mathbf{s}$ raised in previous works [11, 16, 45], in this paper, we mainly use the following 3 importance function: 1) Key Norms (small): $\mathbf{s}=-L_{2}(\mathbf{K}_{g}^{(l)})$; 2) Value Norms: $\mathbf{s}=L_{2}(\mathbf{V}_{g}^{(l)})$; 3) Attention Scores: $\mathbf{s}=\text{matmul}(\mathbf{K}_{g}^{(l)},\mathbf{Q}^{(l)})$. Here $L_{2}$ denotes the L2-norm function and $\mathbf{Q}^{(l)}\in\mathbb{R}^{|\mathbf{X}^{t}|\times(n_{h}\times d_{h})}$ denotes query vector of the text tokens in layer $l$. For our QuickPrefill, we use Key Norms (small) as the default importance function, due to its good performance.
3.3 Overlapping QuickCodec and QuickPrefill
The preceding sections introduced two complementary components: QuickCodec for CPU-based video decoding and QuickPrefill for GPU-based group-wise prefilling. However, sequential execution of these components leads to suboptimal resource utilization; the GPU remains idle during video decoding while the CPU is underutilized during prefill. To address this inefficiency, we propose an overlapped execution scheme that enables concurrent processing across CPU and GPU resources.
Overlapping QuickCodec and QuickPrefill requires a small adaptation to how frames are loaded. Instead of using $c$ cores to load $c$ intervals, we divide $\mathcal{V}$ into $s$ intervals, where $s\gg c$ using Keyframe Intervals($\mathcal{V},s$). We then load the frames from the $s$ intervals using $c$ cores, where the intervals corresponding to earlier blocks of video are loaded first. This approach allows us to use QuickCodec’s fast video decoding, while ensuring that early frames in $\mathcal{V}$ are prioritized, allowing QuickPrefill to begin prefilling groups on GPU. Once the video frames required for the first group are loaded, QuickPrefill immediately begins prefilling. QuickCodec continues to load frames using the CPU in the background. After QuickPrefill finishes prefill for a group, it saves the resulting KV cache vectors and checks if QuickCodec has loaded the frames required to process the next block. If the frames are available, QuickPrefill can begin processing the next group immediately. This establishes a producer-consumer pipeline between CPU video decoding and GPU prefill, where the GPU is only idle if the frames for its next block are not loaded.
The performance benefit of this overlapping execution can be formally analyzed. Let $t_{dec}$ and $t_{prefill}$ denote the total time for decoding and prefilling all video groups, respectively. In the sequential approach, the total execution time is $t_{dec}+t_{prefill}$. However, with our overlapped strategy, the execution time reduces to:
$t_{total}=\max\{t_{dec}+t^{g}_{prefill},t_{prefill}+t^{g}_{dec}\}+\Delta$ (3)
Where $t^{g}_{dec}$ and $t^{g}_{prefill}$ represent the time to decode frames for the first group and prefill the last group, respectively. $\Delta$ is a small amount latency added by QuickCodec parsing the video metadata. As an individual group corresponds to a small number of frames and video tokens, this approach achieves near-optimal overlap between CPU and GPU resources. This results in substantial speedup for hour-scale video processing. We note that some VideoLLMs have additional input preprocessing steps, such as calculating position embeddings or normalization, which we exclude from this analysis [4].
4 Experiments
We evaluate QuickVideo’s performance on practical long video understanding tasks. In section 4.1, we benchmark QuickCodec against existing frameworks. We also examine the limitations of QuickCodec, identifying use-cases where seek-based frameworks (Algorithm 1) have stronger performance. Next, in section 4.2 we evaluate the performance of QuickPrefill on four long video understanding benchmarks and analyze the trade-off between accuracy and efficiency. Lastly, in section 4.3, we demonstrate that our prefill and video decoding stages can be almost entirely overlapped, effectively removing a minute from the time it takes to infer an long-video input.
4.1 QuickCodec Results
Figure 2: Speed comparison of Decord, TorchCodec (with Resize), and QuickCodec on loading hour-long video. We ablate across different levels of parallelization (core counts).
Video Loading Speed. We benchmark the time to load an hour-long 24 FPS 1920x1080p HD video, sampled at 1 FPS, and then resized to 448x448 pixels. The video is an hour segment of a popular movie encoded with default settings in FFmpeg using H.264, the most widely used standard [17]. Sampling frames at $1-2$ FPS is common for VideoLLMs, adopted as a trade-off between computational efficiency and understanding [4]. We resize frames to 448x448 pixels because it is the highest per frame resolution used by most VideoLLMs [48, 7]. We use an AWS m7a.16xlarge instance for our timing. Timings are averaged over five runs, all timings have a $95\%$ confidence interval of at most $0.5$ seconds. We compare QuickVideo against two video decoding frameworks designed for machine learning and integrated into existing long-video understanding inference pipelines:
Decord [12]. Decord is a framework for multimedia loading designed for machine learning applications. Although it is no longer actively maintained, it remains a standard video decoding framework integrated into popular libraries like Hugging Face’s Transformers [39] and, by extension, inference libraries such as vLLM [19].
TorchCodec [27]. TorchCodec is an under-development framework from the PyTorch team. It is designed to offer faster multimedia processing than TorchVision [23]. As it is a work-in-progress library, TorchCodec does not support all the features of more mature frameworks. In particular, TorchCodec does not have built-in support for frame resizing. For this reason, we report timings for the video loading and then a resizing step using TorchVision. TorchCodec is also not designed to use more than 16 cores for video decoding, we find that increasing core count beyond 16 can sometimes decrease TorchDodec’s performance.
As shown in Figure 2, QuickCodec is faster than other libraries across many levels of parallelization. Whereas other frameworks plateau at 16 cores, QuickCodec scales to 64 cores. We highlight the 16 and 32 core cases, as these are the most realistic configurations. Most compute providers configure nodes with between 16 and 32 CPU cores per GPU [15, 3, 24]. Using 16-32 cores, QuickCodec is 2-3 times faster than other libraries at loading an hour-long video, reducing the time to load a video by more than 20 seconds.
Figure 3: Video decoding performance for different video durations with 1 FPS sampling.
Speed across video durations. Our framework requires pre-computing intervals and sufficient keyframes to parallelize over. Therefore, we expect it to perform worse on shorter video durations. We benchmark QuickCodec for different video lengths, ranging from 1 minute to an hour-long video. We use the same hour-long movie video, cut to durations ranging from $1$ to $60$ minutes. We sample at 1 FPS and use $16$ cores for video decoding. All timings are averaged over 5 runs on a AWS m7a.16xlarge instance and have a $95\%$ confidence interval with at most a $0.2$ second margin of error. We find that QuickCodec is faster than other frameworks when loading videos that are more than a minute long (Figure 3). Our framework scales better in long-video loading than other libraries, increasing from 1.7x faster than Decord at loading 10 minute video to 2.1x faster at loading hour-long video. We examine additional limitations of QuickCodec compared to seek-based decoders in Appendix B.
4.2 QuickPrefill Results
In this section, we evaluate the performance of QuickPrefill across four long video understanding benchmarks whose video length ranges from minutes to hours: VideoMME [14], LongVideoBench [40], LVBench [34], and MLVU [46]. All generations are conducted using greedy sampling, and we report results using the lmms-eval framework [44]. All experiments are performed using the Qwen2.5-VL-7B-Instruct model [4] on a single A100 (40GB) GPU with 8 replicas.
Table 1: Effectiveness of different KV cache pruning methods in the group-based prefilling scenario. We use the Key Norms (small) as the default KV cache pruning method for QuickPrefill due to its superior performance and query-agonistic nature.
Group Size KV Pruning $\rho$ VideoMME LongVideoBench LVBench MLVU Avg Performance
#Frames w/o subtitle val test dev
64 Frames
- - 1 62.41 59.69 40.09 63.86 56.51 100.00%
16 Value Norms 0.5 47.63 35.98 30.92 31.38 36.48 64.55%
16 Attention Scores 0.5 58.63 52.95 37.83 59.87 52.32 92.58%
16 Key Norms (small) 0.5 60.56 56.17 37.70 62.34 54.19 95.90%
128 Frames
- - 1 66.41 60.96 42.87 66.86 59.27 100.00%
16 Value Norms 0.5 48.56 37.32 30.73 38.51 38.78 65.42%
16 Attention Scores 0.5 60.96 55.20 39.70 64.36 55.06 92.89%
16 Key Norms (small) 0.5 63.41 58.19 39.57 64.99 56.54 95.39%
256 Frames
- - 1 65.78 61.56 43.90 68.65 59.97 100.00%
16 Value Norms 0.5 48.33 38.89 31.38 37.74 39.08 65.17%
16 Attention Scores 0.5 62.52 57.22 41.96 67.27 57.24 95.45%
16 Key Norms (small) 0.5 64.04 60.21 41.90 66.73 58.22 97.08%
1024 Frames
- - 1 62.00 60.43 42.29 63.48 57.05 100.00%
16 Value Norms 0.5 47.37 33.66 29.18 32.65 35.71 62.60%
16 Attention Scores 0.5 62.22 58.49 42.03 64.45 56.80 99.56%
16 Key Norms (small) 0.5 59.99 61.59 40.80 64.76 56.78 99.53%
Figure 4: Ablation study on the group size and retention ratio. Data from Table 2.
Effectiveness of Different KV Cache Pruning Methods.
We evaluate the impact of various KV cache pruning strategies on model accuracy, as summarized in Table 1. Specifically, we compare several pruning techniques against a baseline with no pruning applied ($\rho=1$). We fix the KV cache retention ratio $\rho$ to 0.5 and set the group size to 16 frames.
Among the evaluated strategies, the Key Norms (small) method achieves the best trade-off between efficiency and accuracy, retaining over 95% of the model’s original performance while reducing KV cache size and computation by half. In the 1024 frames setting, it can even retain more than 98% of the model’s original performance Notably, this method outperforms alternatives that select tokens based on query attention scores. While prior work [11] demonstrated that negative L2 norms of keys correlate strongly with attention scores in text-only LLMs, our results extend this finding to the VideoLLM prefilling setting. This suggests that key norm-based pruning is not only effective during decoding but also applicable in the context of prefilling, underscoring its generalizability and practical value for long-context video understanding.
We further conduct ablation studies to assess the effects of group size and the retention ratio $\rho$ (see Appendix E). As shown in Table 2 and Figure 4, varying the group size has minimal impact on model performance, while increasing the retention ratio $\rho$ consistently improves accuracy, eventually matching the performance of the non-pruned baseline. Smaller group sizes yield lower activation memory, and lower retention ratios result in reduced KV cache memory. These findings offer practical guidance for balancing memory efficiency and model accuracy based on user constraints.
4.3 Latency in end-to-end QuickVideo Inference
Figure 5: Latency from video loading, prefill and LLM decoding in an end-to-end inference setting. We compare a baseline implementation of Qwen2.5-VL [4], Qwen2.5-VL implemented with QuickPrefill and QuickCodec, and lastly, our block overlapped design.
We implement both QuickCodec and QuickPrefill into a Qwen2.5-VL-7B-Instruct [4] inference pipeline. We implement our two versions: loading the entire video input using QuickCodec before prefilling with QuickPrefill, as well as our group overlapped variant. We benchmark latency from the video loading, prefill and LLM decoding stages in our end-to-end inference pipeline. For QuickPrefill, we use the Key Norms (small) as the KV cache pruning method and set $\rho$=0.2, and the number of frames in each group to be 32. We use a 30 minute video (sampled at 1 FPS), as the baseline implementation runs out of memory when using longer video inputs. We use an A100 80GB SXM4 GPU and an AMD EPYC 7513 32-Core CPU. We allocate 16 cores for video processing. For our overlapped implementation, we use 64 video segments ($s$) for our parallelized loading. Our timings are averaged over 10 runs.
Figure 5 shows the latency from video loading, prefill and LLM decoding of all 3 implementations. After QuickCodec reduces the video loading time, we can almost completely overlap video loading and prefill. In the overlapped pipeline, video processing, prefill and LLM decoding completes in only 20.0 seconds, a 49.7 second speedup over the baseline’s 69.7 seconds. Our overlapped implementation has a small startup latency, due to metadata parsing and decoding the frames for the first prefill block. This amounts to 2.8 seconds of video loading that cannot be overlapped with prefill.
5 Discussion and Related Work
GPU support for video decoding. Video decoding can be accelerated by GPU computing. However, due to interframe dependencies, the speedup is not nearly as large as GPU acceleration for AI computations [27]. Furthermore, especially in the case of long video, GPU-based video decoding can result in device memory problems; the hour-long video we use for benchmarking (Section 4.1) is 3600×3×1920×800×1 $3600\times 3\times 1920\times 800\times 1\text{ byte}\approx\boxed{16.6~{}\text{GB}}$ before being resized. This results in a significant portion of GPU resources being allocated to video tensors, and can cause CUDA out-of-memory errors if not handled delicately. For simplicity, most existing inference libraries default to using CPU for video decoding [19, 39]. More sophisticated pipelines, such as NVIDIA’s Cosmos training, use dedicated hardware for handling the video processing [26].
Efficient VideoLLMs Inference. Recent VideoLLMs [21, 20, 7] have demonstrated strong video understanding capabilities. Early models like Video-LLaVA [21] and VideoLLama-2 [9] were limited to around 32 input frames due to constrained training data and unoptimized architectures. More advanced models such as Qwen2.5-VL [4] and InternVideo2.5 [35] can now handle hundreds of frames by adopting architectural innovations including Group Query Attention (GQA) [1], MRoPE [4], and Special Token Merging [7], which reduce KV cache size and enhance temporal reasoning. Nonetheless, the KV cache and activation memory still grow linearly with context length, creating bottlenecks in hour-long video inference. Meanwhile, existing token pruning techniques either address only image-level contexts [36, 6, 28, 42], or optimize for short prefill and long decoding scenarios [11, 45, 41]. In contrast, we target efficient prefill for millions of video tokens, introducing a method that achieves substantial memory savings and speedup with minimal accuracy loss, thereby enabling scalable long video understanding on resource-constrained hardware.
6 Conclusion
We introduced QuickVideo, a framework to accelerate long video understanding. Our framework has three core contributions: QuickCodec: A systems framework for fast video loading, designed for VideoLLM frame sampling. QuickPrefill: An efficient algorithm for prefilling video tokens. Co-design: Lastly, we show that our video loading and prefill algorithm can be almost entirely overlapped, drastically reducing the time latency of these stages during inference. Overall, QuickVideo reduces time to infer a long video input by more than 3×3$\times$3 ×. Our work advances the capabilities for real-time video understanding applications, addressing key efficiency challenges in long video inference.
Appendix A Parallelized Interval Algorithm
Additional video decoding background. The container contains various metadata about packets that we use during our interval parsing algorithm. For locality purposes, modalities such as audio and video are often interleaved in the bit stream $\mathcal{S}$. Therefore, it is important to filter out audio packets when parsing the metadata stream. As packets are not frame-aligned, the pts field does not exactly represent the display time of frame. Also, as packets can be reordered by the decoder, the first or last packets may not correspond to the first and last frames.
Algorithm 3 Calculate Parallelized Intervals
1:procedure Keyframe Intervals($\mathcal{S},c$)
2:     $\mathcal{K},\;pts_{min},\;pts_{max}\leftarrow\textsc{Scan Packets}(\mathcal{S})$triangleright Scan packet metadata.
3:     $\mathcal{J}\leftarrow\{pts_{min},\;pts_{max}\}$triangleright Ordered list of keyframe intervals.
4:     $p\leftarrow\frac{1}{c}({pts_{max}-pts_{min}})$triangleright Evenly spaced intervals in the video.
5:     for $i\in 1,\dots,c-1$ do
6:         $pts_{estimate}\leftarrow(c\times p)+pts_{min}$
7:         $j\leftarrow\textsc{FindInsertionIndex}(\mathcal{K},\;pts_{estimate})$
8:         if $|\mathcal{K}_{j-1}-pts_{estimate}|<|\mathcal{K}_{j}-pts_{estimate}$ then
9:              $\mathcal{J}=\mathcal{J}\cup\{\mathcal{K}_{j-1}\}$
10:         else
11:              $\mathcal{J}=\mathcal{J}\cup\{\mathcal{K}_{j}\}$               return $\mathcal{J}$
12:procedure Scan Packets($\mathcal{S}$) $\triangleright$ Scan bit stream to get timestamps.
13:     $pts_{min}\leftarrow-1$
14:     $pts_{max}\leftarrow\infty$
15:     $\mathcal{K}\leftarrow\emptyset$triangleright Sorted set of keyframe timestamps.
16:     for $p_{i}\in\mathcal{S}$ do
[…]
18:              continue
19:         if $p_{i}.\text{pts}=\texttt{NULL}$ then $\triangleright$ Skip packets do not have pts metadata.
20:              continue
21:         if $p_{i}.\text{pts}<pts_{min}$ then
22:              $_{min}\leftarrow p_{i}.\text{pts}$ pts
23:         if $p_{i}.\text{pts}>pts_{max}$ then
24:              $_{max}\leftarrow p_{i}.\text{pts}$ pts
25:         if $p_{i}.\text{keyframe}=\text{True}$ then
26:              $\mathcal{K}\leftarrow\mathcal{K}\cup\{p_{i}.\text{pts}\}$
27:     return $\mathcal{K},\;pts_{min},\;pts_{max}$
Algorithm 3 computes $c$ intervals that we can parallelize video decoding over. For effective parallelization, it is essential that these intervals are roughly length and keyframe-aligned, such that Algorithm 2 can seek to the start of each interval. Scan Packets parses the metadata of the packet stream to find the location of all keyframes in $\mathcal{S}$, as well as the minimum and maximum pts in $\mathcal{S}$. If the packet does not belong to the video stream or the timestamp is NULL, the packet is skipped.
After finding the locations of keyframes on line 2, Keyframe Intervals computes $c$ intervals as follows: We calculate the length of $\frac{1}{c}$ of the stream, in pts units (line 4). On lines 5-10, we search for the keyframes closest to being $\frac{i}{c}th$ through the video, given by $pts_{estimate}$. FindInsertionIndex uses binary search to find where in the list of keyframes $pts_{estimate}$ would be inserted. After finding the insertion point $j$, the algorithm checks whether the keyframe before or after $j$ is closer to $pts_{estimate}$. The closest keyframe location is added to $\mathcal{J}$, the list of intervals. $\mathcal{J}[0]=pts_{min}$ and $\mathcal{J}[c-1]=pts_{max}$, to ensure that the intervals span the video. $\mathcal{J}[1:c-2]$ are keyframe-aligned and equally spaced. Therefore, $\mathcal{J}$, a list containing $c+1$ values, can be interpreted as $c$ intervals: $\mathcal{J}^{\prime}=\{(\mathcal{J}[i],\mathcal{J}[i+1])\;|\;i\in 0,1,\dots,\;c\}$.
Appendix B Effect of sampling rates on QuickCodec’s efficiency
Figure 6: Video decoding performance for different video durations with 1 FPS sampling.
As QuickCodec does not seek between loading frames, all video frames are decoded during video loading. Conversely, seek-based frameworks skip decoding segments of video if there are large gaps between sampled frames. In Figure 6, we find that our framework has faster video loading when there is a 4 second or less gap between sampled frames. Our library performs best when using VideoLLM sampling rates (1-2 FPS). Currently, our implementation always loads the whole video, and therefore does not benefit significantly from sparse sampling patterns. Our implementation could be adapted to leverage seeking when it detects that the user has sampled with a large gap between frames, closing the performance gap with seek-based libraries [27, 12]. This would make our library more flexible, and eliminate a potential performance sharp edge, where users accidentally use our QuickCodec for sparse sampling. We leave this as a direction for future library improvements.
Appendix C Containers and Video Decoding
A multimedia container file format, like MP4 or MKV, bundles together all the elements required for media playback, including video streams, audio streams, subtitles, images, and metadata [18]. Video streams are compressed into bit streams by codecs . The bit streams are formatted in standards like H.264 [38] and H.265 [31]. A codec consists of two algorithms: a video encoding algorithm that takes in a sequence of frames and outputs a compressed bit stream and a video decoding algorithm that takes the bit stream as input and outputs video frames. We focus video decoding, as it is the required operation before the video can be used as a VideoLLM input.
Appendix D QuickPrefill Efficiency Analysis Details
D.1 Activation Memory Analysis
The activation memory of modern LLM architecture mainly comes from two components of each transformer block: 1) Attention Block and 2) MLP Block. We analyze the potential activation memory usage in formulas in the followings and show that group-based prefilling can effectively reduce the activation memory by $G$ times, where $G$ is the number of groups.
Attention Block
Modern LLMs commonly adopt FlashAttention [10], a memory-efficient attention algorithm that computes exact attention with reduced memory usage by fusing multiple steps and processing attention in blocks. While the naive attention implementation would instantiate the full attention matrix $\bm{A}\in\mathbb{R}^{S\times S}$, FlashAttention avoids this by computing attention block by block. Let $\bm{Q},\bm{K},\bm{V}\in\mathbb{R}^{B\times S\times d_{\text{head}}}$ denote the query, key, and value tensors respectively, with $n_{h}$ heads and $d_{\text{head}}=\frac{d_{\text{model}}}{n_{h}}$. FlashAttention divides the input sequence into blocks of size $B_{c}$ (for keys/values) and $B_{r}$ (for queries) to process attention efficiently within GPU memory constraints. Following [10], the dominant activation memory in FlashAttention comes from storing $\bm{Q}$, $\bm{K}$, $\bm{V}$. The block-based processing means that at any given time, only blocks of the attention matrix of size $B_{r}\times B_{c}$ are materialized in memory. Assume using float16 data type, the total activation memory can be expressed as:
$\mathcal{M}_{\text{attn}}\approx(3B\cdot S\cdot n_{h}\cdot d_{\text{head}}+B\cdot n_{h}\cdot B_{r}\cdot B_{c})\cdot 2\text{ bytes}$ (4)
The first term accounts for storing $\bm{Q}$, $\bm{K}$, and $\bm{V}$ tensors, while the second term accounts for the block of attention matrix being processed. With appropriate block sizes $B_{r}$ and $B_{c}$ (typically set based on GPU memory constraints), the second term remains relatively small. Assuming $B=1$, $S=(|\mathbf{X}^{v}|+|\mathbf{X}^{t}|)\approx 921600$, $d_{\text{model}}=4096$, $n_{h}=8$, $B_{r}=B_{c}=1024$, we compute:
$\mathcal{M}_{\text{attn}}$ $=(3\cdot 1\cdot 921600\cdot 8\cdot 512+1\cdot 8\cdot 1024\cdot 1024)\cdot 2\text{ bytes}$ (5)
$=60,584,722,432\text{ bytes}$ (6)
$\approx\boxed{21.1\text{ GB}}$ (7)
While FlashAttention significantly reduces memory requirements compared to naive attention implementation, this analysis shows it still consumes substantial memory for very long sequences. With group-based prefilling using $G=225$ groups, we can reduce the sequence length $S$ by $G$ times, reducing $\mathcal{M}_{\text{attn}}$ from $21.1$ GB to approximately $\mathbf{0.09}$ GB. This dramatic reduction enables the processing of extremely long sequences that would otherwise be infeasible.
MLP Block
The SwiGLU (Swish-Gated Linear Unit) [29] enhances transformer models through improved gating mechanisms and has been adopted as the default MLP architecture in many popular LLMs including InternVL2.5 and Qwen2.5 series [4, 7]. For input representation $\bm{x}\in\mathbb{R}^{d_{\text{model}}}$, the SwiGLU operation is defined as:
$\text{SwiGLU}(\bm{x})=\bm{W}_{\text{down}}(\text{SiLU}(\bm{W}_{\text{gate}}\bm{x})\odot\bm{W}_{\text{up}}\bm{x})$ (8)
where $\bm{W}_{\text{gate}},\bm{W}_{\text{up}}\in\mathbb{R}^{d_{\text{ff}}\times d_{\text{model}}}$, $\bm{W}_{\text{down}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{ff}}}$, and $\text{SiLU}(x)=x\cdot\sigma(x)$ with $\sigma(x)=\frac{1}{1+e^{-x}}$.
For a batch of sequences, activation memory analysis reveals requirements at each computational step. With batch size $B$, sequence length $S$, hidden dimension $d_{\text{model}}$, intermediate dimension $d_{\text{ff}}$, and data type float16, the total activation memory for a single SwiGLU layer is:
$\mathcal{M}_{\text{act}}=\left(B\cdot S\cdot\left(2d_{\text{model}}+4d_{\text{ff}}\right)\right)\cdot 2\text{ bytes}$ (9)
For a one hour video sampled with 1 FPS ($3600$ frames in total), parameters can be set $B=1$, $S=(|\mathbf{X}^{v}|+|\mathbf{X}^{t}|)\approx 921600$, $d_{\text{model}}=4096$, and $d_{\text{ff}}=14336$:
$\mathcal{M}_{\text{act}}$ $=\left(1\cdot 921600\cdot\left(2\cdot 4096+4\cdot 14336\right)\right)\cdot 2\text{ bytes}$ (10)
$=241,591,910,400\text{ bytes}$ (11)
$\approx\boxed{112.5\text{ GB}}$ (12)
This substantial memory requirement highlights the computational challenges in deploying SwiGLU-based models for high-resolution inputs with extended sequence lengths. However, if we prefill the tokens group by group, we can reduce the $S$ by $G$ times, and thus reduce the activation memory $\mathcal{M}_{\text{act}}$ by $G$ times. Assuming each group contains tokens of 16 frames, then $G=\frac{3600}{16}=225$ and we can reduce $\mathcal{M}_{\text{act}}$ from $112.5$ GB to $\mathbf{0.5}$ GB, which is a substantial improvement.
D.2 KV cache Memory Analysis
When using InternVL2.5-8B [7], with each frame encoded as 256 tokens ($|V|=3{,}600\times 256=921{,}600$), and $|Q|=256$ text tokens, $L=28$ layers, $n_{h}=8$ heads, and $d_{h}=512$, the total memory required to store the KV cache in float16 precision is:
$\text{Memory}=2\times L\times(|\mathbf{X}^{v}|+|\mathbf{X}^{t}|)\times n_{h}\times d_{h}\times 2~{}\text{bytes}\approx\boxed{393.9~{}\text{GB}}.$ (13)
Appendix E Ablation Study on Group Size and Retention Ratio
Table 2: Ablation study of different group sizes and retention ratio $\rho$. We use Key Norms (small) as the KV pruning method here.
Group
Size
$\rho$ VideoMME
LongVideoBench
LVBench
MLVU
Avg Performance
Varying Group Size
- 1 65.78 61.56 43.90 68.65 59.97 100.00%
4 0.5 63.78 60.36 42.61 66.81 58.39 97.36%
8 0.5 64.00 60.88 42.35 66.94 58.54 97.62%
16 0.5 64.04 60.21 41.90 66.73 58.22 97.08%
32 0.5 63.59 59.46 41.51 66.78 57.84 96.44%
64 0.5 63.89 60.51 42.29 66.83 58.38 97.34%
128 0.5 63.56 59.24 42.61 66.97 58.09 96.87%
Varying Retention Ratio $\rho$
16 1 65.78 61.56 43.90 68.65 59.97 100.00%
16 0.1 55.89 53.40 36.02 59.02 51.08 85.18%
16 0.2 59.74 56.47 39.57 61.58 54.34 90.61%
16 0.4 63.22 58.94 41.19 65.75 57.27 95.51%
16 0.6 64.74 60.81 41.90 67.48 58.73 97.93%
16 0.8 65.70 61.41 43.51 68.37 59.75 99.63%
16 0.9 65.85 61.18 43.71 68.70 59.86 99.82%
Appendix F Reproducibility Statement
All machines used for timing experiments are running only essential operating system processes. We report how many runs our results are averaged over, as well $95\%$ confidence intervals constructed using SciPy [33]. We try to use accessible configurations for timings (for example, AWS cloud instances) where possible. All code used for our paper will be open-sourced.
Appendix G Limitations and Broader Impact
Limitations. As it is slow and resource intensive, most VideoLLMs are not trained to use their 1-2 FPS short video sampling rates when using processing long video [4, 7, 48]. Instead, they use very low sampling rates over large time-spans, as we discussed in Section 1. Therefore, VideoLLMs do not (yet) gain a large performance advantage by processing a large number of frames. However, it is clear that a model that has seconds-long gaps between frames can never capture fine-grained temporal and spatial details. Our hope is that making long video understanding (with realistic sampling rates) practical from a systems and algorithm perspective, we will empower the development of such models. Another limitation is that our QuickCodec timings only use H.264 coded video for timings. Although H.264 is the dominant standard, it is not universal.
Broader Impact. As video has become the dominant modality of data, efficient long video understanding has extremely broad implications, both positive and negative. On the positive side, better long video understanding allows us to better interpret our digital landscape. In 2022, 30,000 hours of video were uploaded to YouTube every hour [5]. That number is absolutely much higher today. Without efficient long video understanding systems, we cannot understand our own digital artifacts, due to the scale at which we create them. Furthermore, long video understanding also has extremely compelling use-cases for information accessibility. A video-first internet is difficult to navigate for visually impaired people, with important information potentially only accessible in video format [22]. Efficient, robust long video understanding presents can serve as a backbone for tools for assisting video understanding for the visually impaired. However, efficient long video understanding also has potentially negative effects. As people’s lives are increasingly documented as video and uploaded to the internet, long video understanding models could become a tool for privacy intrusion [13].

### v2

@@
 3 Method
-In this section, we first introduce QuickVideo, mainly consisting of three components:
+In this section, we introduce QuickVideo, which consists of three main components:
 3.1 QuickCodec: Long Video Decoding for VideoLLMs
-Given a bit stream $\mathcal{S}$ for a video $\mathcal{V}=(f_{1},f_{2},\dots,f_{m})$ with frame height $h$, frame width $w$ and a desired degree of concurrency $c$, our goal is to compute ${\bm{F}}$ such that $\forall j\in\{0,1,\dots,|\mathcal{I}|-1\}\;{\bm{F}}_{j}=f_{\mathcal{I}[j]}$ for $\mathcal{I}\subseteq\{1,2,\dots,m-1\}$.
-That is, ${\bm{F}}$ is a packed tensor containing all the frames selected in $\mathcal{I}$.
-We assume that $m$ is known from container metadata or an estimate using $pts_{max}$ and $pts_{min}$.
-The efficiency of our algorithm relies on two observations:
-(1) It is faster to use $c$ cores to decode $c$ short videos than use $c$ cores to sequentially decode $1$ long video.
-Video decoding for human playback focuses on the second case, as humans watch earlier frames while later frames decode.
-However, due to interframe dependencies, sequential video decoding is difficult to parallelize [38].
-Unlike in human playback, VideoLLMs require the entire video input to be loaded upfront.
-Therefore, we can decompose the loading of a long video $\mathcal{V}$ into loading $c$ short videos that span $\mathcal{V}$.
-However, we cannot begin video decoding from an arbitrary frame, only a keyframe.
-Keyframe Intervals (Appendix A) is a subroutine that parses the metadata of $\mathcal{S}$ and computes $c$ approximately equal length intervals, starting and ending on keyframes, that span $\mathcal{V}$.
+Given a bitstream $\mathcal{S}$ for a video $\mathcal{V}=(f_{1},f_{2},\dots,f_{m})$, where each frame has height $h$, width $w$, and the desired degree of concurrency is $c$, our goal is to compute ${\bm{F}}$ such that for all $j\in{0,1,\dots,|\mathcal{I}|-1}$, we have ${\bm{F}}_{j}=f_{\mathcal{I}[j]}$, where $\mathcal{I}\subseteq{1,2,\dots,m-1}$.
+In other words, ${\bm{F}}$ is a packed tensor containing all the frames selected by $\mathcal{I}$.
+We assume $m$ is known from container metadata or can be estimated using $pts_{max}$ and $pts_{min}$.
+The efficiency of our algorithm relies on two key observations:
+(1) It is faster to use $c$ cores to decode $c$ short videos than to use $c$ cores to sequentially decode a single long video.
+Video decoding for human playback focuses on the latter case, as humans watch earlier frames while later frames are decoded.
+However, due to inter-frame dependencies, sequential video decoding is difficult to parallelize [38].
+In contrast, VideoLLMs require the entire video input to be loaded upfront.
+Therefore, we can decompose the loading of a long video $\mathcal{V}$ into loading $c$ short videos that collectively span $\mathcal{V}$.
+However, decoding cannot start at arbitrary frames—it must begin at keyframes.
+The Keyframe Intervals subroutine (Appendix A) parses the metadata of $\mathcal{S}$ and computes $c$ approximately equal-length intervals, starting and ending at keyframes, that cover the entire video.
 We parallelize over these intervals in Algorithm 2.
-(2) Ideally, VideoLLMs sample frames at a short, regular interval, usually 1-2 FPS [4].
-This is less than the gap between keyframes in standard codecs.
-Therefore, seek-based decoding must decode from all keyframes regardless, and many seeks are redundant.
-Our algorithm only requires $1$ seek operation per core, instead of seeks proportional to the number of frames sampled.
+(2) VideoLLMs typically sample frames at a short, regular interval, usually 1–2 FPS [4], which is often smaller than the interval between keyframes in standard codecs.
+Consequently, seek-based decoding must still decode from all keyframes, leading to redundant seeks.
+Our algorithm requires only one seek operation per core, instead of a number of seeks proportional to the number of frames.
 Algorithm 2 QuickCodec
@@
 21:return ${\bm{F}}$
-Algorithm 2 describes the core of our video decoding algorithm.
-The algorithm begins by using metadata to calculate $c$ keyframe-aligned intervals span the video (line 1).
-Lines 2-5 initialize a block of shared memory ${\bm{F}}$ and compute a dictionary $M$ that maps indices of selected frames to unique memory offsets in ${\bm{F}}$.
-We then decode the long video in $c$ parallel intervals (lines 6-19).
-Video decoding begins by seeking the the start of the interval $pts_{start}$, which is guaranteed to be a keyframe (line 7).
-We enqueue packet to decode (lines 17-18) until the video decoder yields frames to process (line 9).
-If the timestamp of the dequeued frame is greater than or equal to the endpoint of the interval $pts_{end}$, parallel processing ends (lines 11-12 and 19).
-As the intervals in $\mathcal{J}$ span $\mathcal{S}$, $pts_{min}$ and $pts_{max}$ are given by the least and greatest values in $\mathcal{J}$, respectively.
-Therefore, we use can Equation 1 to calculate the index $i$ of $f$ (line 13).
-Lastly, we save $f$ to ${\bm{F}}$ if $f$ is a selected frame (lines 14-16).
-As decoding from a keyframe yields all frames with greater $pts$ and $\mathcal{D}$ outputs frames in $pts$ order, when the parallelized loop exits (line 19), all selected frames with $pts$ in the interval $[pts_{start},pts_{end})$ will have been output by $\mathcal{D}$ and saved to ${\bm{F}}$.
-Therefore, as $\mathcal{J}$ spans $\mathcal{S}$, when the algorithm returns, ${\bm{F}}$ will contain all selected frames.
-3.2 QuickPrefill: Efficient Group-based Pre-filling for VideoLLM
-After decoding the video bit stream into packed tensors, they are then fed into the VideoLLM for inference.
-However, LLM generation with a long context is a well-known challenge due to its intensive memory usage and costly computation.
-To mitigate this issue, we introduce QuickPrefill, a grouped prefilling and KV cache pruning method that both accelerates the speed and reduces the memory requirement significantly.
-Group-based Prefill.
-Let $V={\mathbf{v}_{1},\mathbf{v}2,\ldots,\mathbf{v}{|\mathbf{X}^{v}|}}$ denote the sequence of video tokens, where $|\mathbf{X}^{v}|$ is the total number of tokens and each token $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional vector.
-To reduce memory overhead during prefilling, we adopt a group-based strategy by partitioning the video token sequence into $G$ disjoint groups: $\mathbf{V}={\mathbf{V}_{1},\ldots,\mathbf{V}_{G}}$, where each group $\mathbf{V}_{g}$ contains approximately $N_{g}=\frac{|\mathbf{X}^{v}|}{G}$ tokens.
-Instead of processing the entire sequence at once, we sequentially prefill each group and store its corresponding key-value (KV) cache as $\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}$ for each transformer layer $l$.
-This strategy significantly reduces the peak activation memory usage by a factor of $G$ and remains effective even when used in conjunction with efficient attention mechanisms such as FlashAttention.
-Empirically, it enables hour-long video understanding while keeping GPU memory usage within practical limits (e.g., reducing more than 100 GB of memory overhead; see subsection D.2).
-Group-based KV Cache Pruning.
-While group-based prefill can effectively reduce the peak activation memory, the largest memory bottleneck, KV cache memory is still not resolved.
-Therefore, when processing each group of video tokens, instead of saving the all KV cache vectors, we will prune unimportant KV cache to maintain an $\rho\in(0,1]$ retention ratio and thus reduce the KV cache memory usage by a factor of $\frac{1}{\rho}$.
-The pruning decision is based on an importance score function $\mathbf{s}$ to produce an ordered list where we select the top-k KV cache until the retention ratio is reached:
-$\tilde{\mathbf{K}}_{g}^{(l)}=\mathbf{K}_{g}^{(l)}[I_{g}^{(l)}],\text{ }\tilde{\mathbf{V}}_{g}^{(l)}=\mathbf{V}_{g}^{(l)}[I_{g}^{(l)}],\text{ where }I_{g}^{(l)}=\operatorname{TopK}\left(\mathbf{s}(\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}),\;k=\rho\cdot N_{g}\right)$ (2)
-where $\operatorname{TopK}$ defines a function that returns of the indices the top-k largest inputs.
-There are multiple heuristic importance score function $\mathbf{s}$ raised in previous works [11, 16, 45], in this paper, we mainly use the following 3 importance function: 1) Key Norms (small): $\mathbf{s}=-L_{2}(\mathbf{K}_{g}^{(l)})$; 2) Value Norms: $\mathbf{s}=L_{2}(\mathbf{V}_{g}^{(l)})$; 3) Attention Scores: $\mathbf{s}=\text{matmul}(\mathbf{K}_{g}^{(l)},\mathbf{Q}^{(l)})$.
-Here $L_{2}$ denotes the L2-norm function and $\mathbf{Q}^{(l)}\in\mathbb{R}^{|\mathbf{X}^{t}|\times(n_{h}\times d_{h})}$ denotes query vector of the text tokens in layer $l$.
-For our QuickPrefill, we use Key Norms (small) as the default importance function, due to its good performance.
+Algorithm 2 presents the core of our video decoding algorithm.
+The algorithm begins by using metadata to compute $c$ keyframe-aligned intervals that span the video (line 1).
+Lines 2–5 initialize a shared memory block ${\bm{F}}$ and compute a dictionary $M$ that maps indices of selected frames to unique memory offsets in ${\bm{F}}$.
+We then decode the long video in $c$ parallel intervals (lines 6–19).
+Video decoding starts by seeking to the start of each interval $pts_{start}$, which is guaranteed to be a keyframe (line 7).
+Packets are enqueued for decoding (lines 17–18) until the decoder yields frames for processing (line 9).
+If the timestamp of a dequeued frame is greater than or equal to the interval endpoint $pts_{end}$, parallel processing ends (lines 11–12, 19).
+As the intervals in $\mathcal{J}$ span $\mathcal{S}$, $pts_{min}$ and $pts_{max}$ are given by the smallest and largest values in $\mathcal{J}$, respectively.
+We use Equation 1 to compute the index $i$ of $f$ (line 13).
+Finally, we save $f$ to ${\bm{F}}$ if $f$ is a selected frame (lines 14–16).
+Because decoding from a keyframe yields all frames with greater $pts$ values, and $\mathcal{D}$ outputs frames in $pts$ order, when the parallelized loop exits (line 19), all selected frames with $pts$ in the interval $[pts_{start},pts_{end})$ will have been output by $\mathcal{D}$ and saved to ${\bm{F}}$.
+Thus, as $\mathcal{J}$ spans $\mathcal{S}$, when the algorithm returns, ${\bm{F}}$ will contain all selected frames.
+3.2 QuickPrefill: Efficient Group-Based Prefilling for VideoLLMs
+After decoding the video bitstream into packed tensors, they are fed into the VideoLLM for inference.
+However, LLM generation with long contexts is a well-known challenge due to high memory usage and computational cost.
+To address this, we introduce QuickPrefill, a grouped prefilling and KV cache pruning method that accelerates processing and significantly reduces memory requirements.
+Group-Based Prefill Let $\mathbf{X}^{v}={\mathbf{x}_{1}^{v},\ldots,\mathbf{x}_{|\mathbf{X}^{v}|}^{v}}$ denote the sequence of video tokens, where $|\mathbf{X}^{v}|$ is the total number of tokens, and each token $\mathbf{x}_{i}^{v}\in\mathbb{R}^{d}$ is a $d$-dimensional vector.
+To reduce memory overhead during prefilling, we partition the video token sequence into $G$ disjoint groups: $\mathbf{X}^{v}={\mathbf{X}^{v}_{1},\ldots,\mathbf{X}^{v}_{G}}$, where each group $\mathbf{X}_{g}^{v}$ contains approximately $N_{g}=\frac{|\mathbf{X}^{v}|}{G}$ tokens.
+Instead of processing the entire sequence at once, we sequentially prefill each group and store the corresponding key-value (KV) cache as $\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}$ for each transformer layer $l$.
+This strategy significantly reduces peak activation memory usage by a factor of $G$ and remains effective even when combined with efficient attention mechanisms such as FlashAttention.
+Empirically, it enables hour-long video understanding while keeping GPU memory usage within practical limits (e.g., reducing memory overhead by over 100 GB; see subsection D.2).
+Group-Based KV Cache Pruning While group-based prefill reduces peak activation memory, the KV cache memory remains a major bottleneck.
+To address this, we prune unimportant KV cache vectors when processing each group, maintaining a retention ratio $\rho\in(0,1]$.
+This reduces KV cache memory usage by a factor of $\frac{1}{\rho}$.
+The pruning decision is based on an importance score function $\mathbf{s}$, which produces an ordered list of KV entries. We select the top-$k$ KV entries until the retention ratio is satisfied:
+$\tilde{\mathbf{K}}_{g}^{(l)}=\mathbf{K}_{g}^{(l)}[I_{g}^{(l)}],\quad\tilde{\mathbf{V}}_{g}^{(l)}=\mathbf{V}_{g}^{(l)}[I_{g}^{(l)}],\quad\text{where }I_{g}^{(l)}=\operatorname{TopK}\left(\mathbf{s}(\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}),;k=\rho\cdot N_{g}\right)$ (2)
+where $\operatorname{TopK}$ returns the indices of the top-$k$ entries.
+We consider several heuristic importance functions $\mathbf{s}$ from prior works [11, 16, 45].
+In this paper, we primarily use the following three: 1) Key Norms (small): $\mathbf{s}=-L_{2}(\mathbf{K}_{g}^{(l)})$; 2) Value Norms: $\mathbf{s}=L_{2}(\mathbf{V}_{g}^{(l)})$; 3) Attention Scores: $\mathbf{s}=\text{matmul}(\mathbf{K}_{g}^{(l)},\mathbf{Q}^{(l)})$.
+Here, $L_{2}$ denotes the L2-norm function, and $\mathbf{Q}^{(l)}\in\mathbb{R}^{|\mathbf{X}^{t}|\times(n_{h}\times d_{h})}$ is the query vector of text tokens in layer $l$.
+We adopt Key Norms (small) as the default importance function in QuickPrefill due to its strong performance.
 3.3 Overlapping QuickCodec and QuickPrefill
 The preceding sections introduced two complementary components: QuickCodec for CPU-based video decoding and QuickPrefill for GPU-based group-wise prefilling.
-However, sequential execution of these components leads to suboptimal resource utilization; the GPU remains idle during video decoding while the CPU is underutilized during prefill.
+However, running these components sequentially underutilizes resources—GPUs remain idle during video decoding, and CPUs are underutilized during prefilling.
 To address this inefficiency, we propose an overlapped execution scheme that enables concurrent processing across CPU and GPU resources.
-Overlapping QuickCodec and QuickPrefill requires a small adaptation to how frames are loaded.
-Instead of using $c$ cores to load $c$ intervals, we divide $\mathcal{V}$ into $s$ intervals, where $s\gg c$ using Keyframe Intervals($\mathcal{V},s$).
-We then load the frames from the $s$ intervals using $c$ cores, where the intervals corresponding to earlier blocks of video are loaded first.
-This approach allows us to use QuickCodec’s fast video decoding, while ensuring that early frames in $\mathcal{V}$ are prioritized, allowing QuickPrefill to begin prefilling groups on GPU.
-Once the video frames required for the first group are loaded, QuickPrefill immediately begins prefilling.
-QuickCodec continues to load frames using the CPU in the background.
-After QuickPrefill finishes prefill for a group, it saves the resulting KV cache vectors and checks if QuickCodec has loaded the frames required to process the next block.
-If the frames are available, QuickPrefill can begin processing the next group immediately.
-This establishes a producer-consumer pipeline between CPU video decoding and GPU prefill, where the GPU is only idle if the frames for its next block are not loaded.
-The performance benefit of this overlapping execution can be formally analyzed.
+To achieve this, we slightly adapt frame loading: Instead of using $c$ cores to load $c$ intervals, we divide $\mathcal{V}$ into $s$ intervals, where $s\gg c$, using Keyframe Intervals($\mathcal{V},s$).
+We then load the frames from the $s$ intervals using $c$ cores, prioritizing earlier intervals so that frames corresponding to the first blocks of video are available sooner.
+This allows us to exploit QuickCodec’s fast video decoding while ensuring that early frames in $\mathcal{V}$ are prioritized for QuickPrefill.
+Once the frames required for the first group are loaded, QuickPrefill begins processing immediately, while QuickCodec continues decoding subsequent frames in the background.
+After QuickPrefill finishes processing a group, it stores the resulting KV cache and checks whether QuickCodec has loaded the frames required for the next group.
+If so, QuickPrefill starts processing the next group immediately.
+This design forms a producer-consumer pipeline between CPU decoding and GPU prefilling, ensuring the GPU is only idle if it is waiting for the CPU to finish decoding the next set of frames.
+The performance improvement of this overlap strategy can be formalized.
 Let $t_{dec}$ and $t_{prefill}$ denote the total time for decoding and prefilling all video groups, respectively.
 In the sequential approach, the total execution time is $t_{dec}+t_{prefill}$.
-However, with our overlapped strategy, the execution time reduces to:
-$t_{total}=\max\{t_{dec}+t^{g}_{prefill},t_{prefill}+t^{g}_{dec}\}+\Delta$ (3)
-Where $t^{g}_{dec}$ and $t^{g}_{prefill}$ represent the time to decode frames for the first group and prefill the last group, respectively.
-$\Delta$ is a small amount latency added by QuickCodec parsing the video metadata.
-As an individual group corresponds to a small number of frames and video tokens, this approach achieves near-optimal overlap between CPU and GPU resources.
-This results in substantial speedup for hour-scale video processing.
-We note that some VideoLLMs have additional input preprocessing steps, such as calculating position embeddings or normalization, which we exclude from this analysis [4].
+In contrast, with our overlap strategy, the execution time is:
+$t_{total}=\max{t_{dec}+t^{g}_{prefill},;t_{prefill}+t^{g}_{dec}}+\Delta$ (3)
+where $t^{g}_{dec}$ and $t^{g}_{prefill}$ represent the time to decode frames for the first group and to prefill the last group, respectively, and $\Delta$ is a small latency introduced by QuickCodec’s metadata parsing.
+Since each group contains a small number of frames and tokens, this strategy achieves near-optimal overlap between CPU and GPU resources, resulting in substantial speedup for hour-long video processing.
+Note that some VideoLLMs include additional preprocessing steps (e.g., position embedding calculation or normalization), which we do not include in this analysis [4].
 4 Experiments
@@
 In section 4.1, we benchmark QuickCodec against existing frameworks.
-We also examine the limitations of QuickCodec, identifying use-cases where seek-based frameworks (Algorithm 1) have stronger performance.
-Next, in section 4.2 we evaluate the performance of QuickPrefill on four long video understanding benchmarks and analyze the trade-off between accuracy and efficiency.
-Lastly, in section 4.3, we demonstrate that our prefill and video decoding stages can be almost entirely overlapped, effectively removing a minute from the time it takes to infer an long-video input.
+We also examine the limitations of QuickCodec, identifying use cases where seek-based frameworks (Algorithm 1) outperform our method.
+Next, in section 4.2, we evaluate the performance of QuickPrefill across four long video understanding benchmarks, analyzing the trade-off between accuracy and efficiency.
+Finally, in section 4.3, we demonstrate that the prefill and video decoding stages can be almost entirely overlapped, effectively reducing end-to-end inference time by nearly a minute for long video inputs.
 4.1 QuickCodec Results
-Figure 2: Speed comparison of Decord, TorchCodec (with Resize), and QuickCodec on loading hour-long video. We ablate across different levels of parallelization (core counts).
+Figure 2: Speed comparison of Decord, TorchCodec (with Resize), and QuickCodec when loading hour-long videos. We ablate across different levels of parallelization (core counts).
 Video Loading Speed.
-We benchmark the time to load an hour-long 24 FPS 1920x1080p HD video, sampled at 1 FPS, and then resized to 448x448 pixels.
-The video is an hour segment of a popular movie encoded with default settings in FFmpeg using H.264, the most widely used standard [17].
-Sampling frames at $1-2$ FPS is common for VideoLLMs, adopted as a trade-off between computational efficiency and understanding [4].
-We resize frames to 448x448 pixels because it is the highest per frame resolution used by most VideoLLMs [48, 7].
-We use an AWS m7a.16xlarge instance for our timing.
-Timings are averaged over five runs, all timings have a $95\%$ confidence interval of at most $0.5$ seconds.
-We compare QuickVideo against two video decoding frameworks designed for machine learning and integrated into existing long-video understanding inference pipelines:
-Decord [12].
-Decord is a framework for multimedia loading designed for machine learning applications.
-Although it is no longer actively maintained, it remains a standard video decoding framework integrated into popular libraries like Hugging Face’s Transformers [39] and, by extension, inference libraries such as vLLM [19].
-TorchCodec [27].
-TorchCodec is an under-development framework from the PyTorch team.
-It is designed to offer faster multimedia processing than TorchVision [23].
-As it is a work-in-progress library, TorchCodec does not support all the features of more mature frameworks.
-In particular, TorchCodec does not have built-in support for frame resizing.
-For this reason, we report timings for the video loading and then a resizing step using TorchVision.
-TorchCodec is also not designed to use more than 16 cores for video decoding, we find that increasing core count beyond 16 can sometimes decrease TorchDodec’s performance.
-As shown in Figure 2, QuickCodec is faster than other libraries across many levels of parallelization.
-Whereas other frameworks plateau at 16 cores, QuickCodec scales to 64 cores.
-We highlight the 16 and 32 core cases, as these are the most realistic configurations.
-Most compute providers configure nodes with between 16 and 32 CPU cores per GPU [15, 3, 24].
-Using 16-32 cores, QuickCodec is 2-3 times faster than other libraries at loading an hour-long video, reducing the time to load a video by more than 20 seconds.
-Figure 3: Video decoding performance for different video durations with 1 FPS sampling.
+We benchmark the time required to load an hour-long 24 FPS 1920x1080p HD video, sampled at 1 FPS and resized to 448x448 pixels.
+The video is a one-hour segment of a popular movie encoded with default FFmpeg settings using H.264, the most widely used codec [17].
+Sampling frames at 1–2 FPS is a standard practice in VideoLLMs, balancing computational efficiency with task performance [4].
+We resize frames to 448x448 pixels, which matches the maximum per-frame resolution used in most VideoLLMs [48, 7].
+All experiments are conducted on an AWS m7a.16xlarge instance.
+Each timing result is averaged over five runs, with a 95% confidence interval no greater than 0.5 seconds.
+We compare QuickVideo against two widely used video decoding frameworks:
+Decord[12]: A multimedia loading framework designed for machine learning applications.
+While no longer actively maintained, Decord remains integrated into popular libraries like Hugging Face’s Transformers[39] and, by extension, inference frameworks such as vLLM [19].
+TorchCodec[27]: A work-in-progress library from the PyTorch team designed to offer faster multimedia processing than TorchVision[23].
+TorchCodec lacks some features of mature frameworks, such as built-in support for frame resizing.
+Thus, we report timings that combine TorchCodec loading with a resizing step via TorchVision.
+TorchCodec is not optimized for decoding with more than 16 cores; we observe that increasing core count beyond 16 can even degrade performance.
+As shown in Figure 2, QuickCodec outperforms other libraries across varying core counts.
+While other frameworks plateau at 16 cores, QuickCodec scales up to 64 cores.
+We highlight the 16- and 32-core cases as these are the most common configurations in practical deployments; most compute providers allocate between 16 and 32 CPU cores per GPU [15, 3, 24].
+At 16–32 cores, QuickCodec is 2–3$\times$ faster than other libraries when loading an hour-long video, reducing video loading time by over 20 seconds.
+Figure 3: Video decoding performance across different video durations (1 FPS sampling).
 Speed Across Video Durations.
-Our framework requires pre-computing intervals and sufficient keyframes to parallelize over.
-Therefore, we expect it to perform worse on shorter video durations.
-We benchmark QuickCodec for different video lengths, ranging from 1 minute to an hour-long video.
-We use the same hour-long movie video, cut to durations ranging from $1$ to $60$ minutes.
-We sample at 1 FPS and use $16$ cores for video decoding.
-All timings are averaged over 5 runs on a AWS m7a.16xlarge instance and have a $95\%$ confidence interval with at most a $0.2$ second margin of error.
-We find that QuickCodec is faster than other frameworks when loading videos that are more than a minute long (Figure 3).
-Our framework scales better in long-video loading than other libraries, increasing from 1.7x faster than Decord at loading 10 minute video to 2.1x faster at loading hour-long video.
-We examine additional limitations of QuickCodec compared to seek-based decoders in Appendix B.
+Our framework relies on pre-computing intervals and sufficient keyframes for parallelization.
+Therefore, we expect reduced benefits for shorter videos.
+We benchmark QuickCodec on videos of varying lengths, from 1 minute to 1 hour, using the same source video (an hour-long movie) cut to different durations.
+All tests use 1 FPS sampling and 16 cores for video decoding.
+Results are averaged over five runs on an AWS m7a.16xlarge instance, with a 95% confidence interval of at most 0.2 seconds.
+We find that QuickCodec is consistently faster than other frameworks for videos longer than 1 minute (Figure 3).
+Its advantage grows with video length—QuickCodec is 1.7$\times$ faster than Decord for a 10-minute video and 2.1$\times$ faster for a 1-hour video.
+We further discuss scenarios where seek-based decoders outperform QuickCodec in Appendix B.
 4.2 QuickPrefill Results
-In this section, we evaluate the performance of QuickPrefill across four long video understanding benchmarks whose video length ranges from minutes to hours: VideoMME [14], LongVideoBench [40], LVBench [34], and MLVU [46].
-All generations are conducted using greedy sampling, and we report results using the lmms-eval framework [44].
-All experiments are performed using the Qwen2.5-VL-7B-Instruct model [4] on a single A100 (40GB) GPU with 8 replicas.
+We evaluate QuickPrefill on four long video understanding benchmarks, with videos ranging from minutes to hours: VideoMME [14], LongVideoBench [40], LVBench [34], and MLVU [46].
+All generations use greedy sampling, and results are reported via the lmms-eval framework [44].
+Experiments are run on the Qwen2.5-VL-7B-Instruct model [4] using a single A100 (40GB) GPU with 8 replicas.
 Table 1: Effectiveness of different KV cache pruning methods in the group-based prefilling scenario.
@@
 16 Key Norms (small) 0.5 59.99 61.59 40.80 64.76 56.78 99.53%
-Figure 4: Ablation study on the group size and retention ratio. Data from Table 2.
+Figure 4: Ablation study on group size and retention ratio. Data from Table 2.
 Effectiveness of Different KV Cache Pruning Methods.
 We evaluate the impact of various KV cache pruning strategies on model accuracy, as summarized in Table 1.
-Specifically, we compare several pruning techniques against a baseline with no pruning applied ($\rho=1$).
-We fix the KV cache retention ratio $\rho$ to 0.5 and set the group size to 16 frames.
-Among the evaluated strategies, the Key Norms (small) method achieves the best trade-off between efficiency and accuracy, retaining over 95% of the model’s original performance while reducing KV cache size and computation by half.
-In the 1024 frames setting, it can even retain more than 98% of the model’s original performance Notably, this method outperforms alternatives that select tokens based on query attention scores.
-While prior work [11] demonstrated that negative L2 norms of keys correlate strongly with attention scores in text-only LLMs, our results extend this finding to the VideoLLM prefilling setting.
-This suggests that key norm-based pruning is not only effective during decoding but also applicable in the context of prefilling, underscoring its generalizability and practical value for long-context video understanding.
-We further conduct ablation studies to assess the effects of group size and the retention ratio $\rho$ (see Appendix E).
-As shown in Table 2 and Figure 4, varying the group size has minimal impact on model performance, while increasing the retention ratio $\rho$ consistently improves accuracy, eventually matching the performance of the non-pruned baseline.
-Smaller group sizes yield lower activation memory, and lower retention ratios result in reduced KV cache memory.
-These findings offer practical guidance for balancing memory efficiency and model accuracy based on user constraints.
+We compare several pruning techniques against a no-pruning baseline ($\rho=1$), fixing the retention ratio $\rho$ at 0.5 and the group size at 16 frames.
+The Key Norms (small) method achieves the best balance between efficiency and accuracy, retaining over 95% of the model’s original performance while halving the KV cache size and computation.
+In the 1024-frame setting, it retains over 98% of the original performance.
+Notably, this method outperforms query-attention-based token selection strategies.
+While prior work [11] has shown that negative L2 norms of keys correlate strongly with attention scores in text-only LLMs, our results extend this finding to VideoLLM prefilling, highlighting the generalizability and practical utility of key norm-based pruning.
+We also conduct ablation studies on group size and retention ratio $\rho$ (see Appendix E).
+As shown in Table 2 and Figure 4, group size has minimal impact on model performance, while increasing $\rho$ consistently improves accuracy, approaching the no-pruning baseline.
+Smaller group sizes reduce activation memory, while lower $\rho$ values reduce KV cache memory.
+These findings provide practical guidance for balancing memory efficiency and model accuracy based on system constraints.
 4.3 Latency in End-to-End QuickVideo Inference
-Figure 5: Latency from video loading, prefill and LLM decoding in an end-to-end inference setting.
-We compare a baseline implementation of Qwen2.5-VL [4], Qwen2.5-VL implemented with QuickPrefill and QuickCodec, and lastly, our block overlapped design.
-We implement both QuickCodec and QuickPrefill into a Qwen2.5-VL-7B-Instruct [4] inference pipeline.
-We implement our two versions: loading the entire video input using QuickCodec before prefilling with QuickPrefill, as well as our group overlapped variant.
-We benchmark latency from the video loading, prefill and LLM decoding stages in our end-to-end inference pipeline.
-For QuickPrefill, we use the Key Norms (small) as the KV cache pruning method and set $\rho$=0.2, and the number of frames in each group to be 32.
-We use a 30 minute video (sampled at 1 FPS), as the baseline implementation runs out of memory when using longer video inputs.
-We use an A100 80GB SXM4 GPU and an AMD EPYC 7513 32-Core CPU.
-We allocate 16 cores for video processing.
-For our overlapped implementation, we use 64 video segments ($s$) for our parallelized loading.
-Our timings are averaged over 10 runs.
-Figure 5 shows the latency from video loading, prefill and LLM decoding of all 3 implementations.
-After QuickCodec reduces the video loading time, we can almost completely overlap video loading and prefill.
-In the overlapped pipeline, video processing, prefill and LLM decoding completes in only 20.0 seconds, a 49.7 second speedup over the baseline’s 69.7 seconds.
-Our overlapped implementation has a small startup latency, due to metadata parsing and decoding the frames for the first prefill block.
-This amounts to 2.8 seconds of video loading that cannot be overlapped with prefill.
+Figure 5: Latency breakdown for video loading, prefill, and LLM decoding in end-to-end inference.
+We compare a baseline Qwen2.5-VL [4] implementation, the same model with QuickPrefill and QuickCodec, and our block-overlapped design.
+We integrate QuickCodec and QuickPrefill into a Qwen2.5-VL-7B-Instruct [4] inference pipeline.
+We evaluate two configurations: (1) loading the entire video with QuickCodec followed by QuickPrefill, and (2) our group-overlapped design.
+Latency for video loading, prefill, and LLM decoding is benchmarked in an end-to-end pipeline.
+For QuickPrefill, we use the Key Norms (small) pruning method with $\rho=0.2$ and set the group size to 32 frames.
+We use a 30-minute video (sampled at 1 FPS) as the baseline implementation runs out of memory with longer videos.
+Experiments are conducted on an A100 80GB SXM4 GPU with an AMD EPYC 7513 32-core CPU, allocating 16 cores for video processing.
+For the overlapped implementation, we use 64 intervals ($s$ in section 3.3) for parallelized loading.
+All timings are averaged over 10 runs.
+Figure 5 presents latency breakdowns for all three implementations.
+After applying QuickCodec, we significantly reduce video loading time.
+By overlapping video loading and prefill, we achieve near-complete overlap of the two stages.
+The block-overlapped design completes video processing, prefill, and LLM decoding in 20.0 seconds, a 49.7-second speedup over the baseline’s 69.7 seconds.
+Our overlapped pipeline introduces a small startup latency—2.8 seconds—for metadata parsing and decoding frames for the first prefill block, which cannot be overlapped.
 5 Discussion and Related Work
@@
 7:         $j\leftarrow\textsc{FindInsertionIndex}(\mathcal{K},\;pts_{estimate})$
-8:         if $|\mathcal{K}_{j-1}-pts_{estimate}|<|\mathcal{K}_{j}-pts_{estimate}$ then
+8:         if $|\mathcal{K}_{j-1}-pts_{estimate}|<|\mathcal{K}_{j}-pts_{estimate}|$ then
 9:              $\mathcal{J}=\mathcal{J}\cup\{\mathcal{K}_{j-1}\}$
@@
 $\mathcal{J}[0]=pts_{min}$ and $\mathcal{J}[c-1]=pts_{max}$, to ensure that the intervals span the video.
-$\mathcal{J}[1:c-2]$ are keyframe-aligned and equally spaced.
+$\mathcal{J}[1],\mathcal{J}[2],\dots\mathcal{J}[{c-2}]$ are keyframe-aligned and equally spaced.
 Therefore, $\mathcal{J}$, a list containing $c+1$ values, can be interpreted as $c$ intervals: $\mathcal{J}^{\prime}=\{(\mathcal{J}[i],\mathcal{J}[{i+1}])\;|\;i\in 0,1,\dots,\;c\}$.